The height of a box is 4 inches greater than its width, the length is 1 inch less than the width. If the box has a volume of 12 cubic inches, what are the dimensions of the box?
Answer by stanbon(75887)

The height of a box is 4 inches greater than its width, the length is 1 inch less than the width. If the box has a volume of 12 cubic inches, what are the dimensions of the box?
-------
Let the width be "x"
Height = x+4
length = x-1
-----
Volume = lwh
12 = x(x-1)(x+4)
x(x^2+3x-4) = 12
x^3 + 3x^2 -4x - 12 = 0
------------------------
Usable solution:
x = 2.3 inches (width)
-------
x+4 = 6.3 inches (height)
-------
x-1 = 1.3 inches (length)
=============================
Cheers,
Stan H.
===============